OpenStudy (anonymous):

How do u find the area of a rectangle that is (x-1) (x+1) equals 150

OpenStudy (anonymous):

x^2=151, is it a rectangle?

OpenStudy (asnaseer):

area of 150 is equal to the product of (x-1) and (x+1) and I assume you need to solve for x? If that is the case, then we have:\[(x-1)(x+1)=150\]\[x^2-1=150\]\[x^2=150+1=151\]therefore:\[x=\sqrt{151}\]as @Yifan12879 indicated above.